What is the mole fraction of NaCl if 5 moles of NaCl is dissolved in 1 L of water?
Answer #1

Volume of water is 1 L = 1000 mL

Since density of water is 1 g/mL,

Mass of water = 1000 g

Molar mass of H2O,

MM = 2*MM(H) + 1*MM(O)

= 2*1.008 + 1*16.0

= 18.016 g/mol

mass(H2O)= 1000 g

use:

number of mol of H2O,

n = mass of H2O/molar mass of H2O

=(1*10^3 g)/(18.02 g/mol)

= 55.51 mol

Now use:

Mole fraction of NaCl = mol of NaCl / Total number of mol

= 5 / (55.51 + 5)

= 0.0826

Answer: 0.0826